### v2

# Entanglement Rényi Negativity of Interacting Fermions from Quantum Monte Carlo Simulations

Abstract
Many-body entanglement unveils additional aspects of quantum matter and offers insights into strongly correlated physics. While ground-state entanglement has received much attention in the past decade, the study of mixed-state quantum entanglement using negativity in interacting fermionic systems remains largely unexplored. We demonstrate that the partially transposed density matrix of interacting fermions, similar to their reduced density matrix, can be expressed as a weighted sum of Gaussian states describing free fermions, enabling the calculation of rank-$n$ Rényi negativity within the determinant quantum Monte Carlo framework. We conduct the first calculation of rank-two Rényi negativity for the half-filled Hubbard model and the spinless $t$-$V$ model and find that the area law coefficient of the Rényi negativity has a logarithmic finite-size scaling at the finite-temperature transition point. Our work contributes to the calculation of entanglement and sets the stage for future studies on quantum entanglement in various fermionic many-body mixed states.
Introduction.—The characterization of emerging quantum many-body phenomena is multifaceted. Traditionally, physicists have relied on local measurements based on linear response to investigate matter. In recent decades, the utilization of quantum entanglement, a fundamental concept in quantum physics and a powerful tool in quantum information, has become pivotal in unveiling the additional aspects of quantum matter, including the identification of exotic phases and quantum criticality [1, 2, 3]. A prominent example is the entanglement entropy (EE) used in bipartite ground-state entanglement studies [4], where various corrections to the area law [5, 6, 7] have been employed to classify quantum phases. These include logarithmic corrections in the leading area-law term for 1D critical systems [8, 9] and Fermi surfaces in generic dimensions [10, 11, 12], subleading logarithmic terms for corner contributions to 2D critical systems [13, 14] and Goldstone modes in symmetry-breaking phases [15, 16, 17, 18], and the topological EE for non-local orders [19, 20, 21, 22, 23].
However, EE is not a faithful mixed-state entanglement measurement due to its incompetence in distinguishing quantum entanglement from classical correlation. Thus, many entanglement measurements for mixed states have been proposed [24, 25], including the entanglement negativity [26, 27, 28, 29] (referred to as “negativity” henceforth for brevity), which was designed base on positive partial transpose criteria for the separability of density matrices [30, 31]. The evaluation of negativity hinges on the partial transpose of the given density matrix and can be carried out straightforwardly through basic matrix manipulations without invoking any optimization. Hence, negativity has been employed to examine entanglement in finite-temperature Gibbs states or tripartite ground states in various systems, spanning from one-dimensional conformal field theory [32, 33, 34, 35], bosonic systems [36, 37, 38, 39, 40], spin systems [41, 42, 43, 44, 45, 46, 47], to topologically ordered phases [48, 49, 50, 51, 52, 53].
In the case of fermionic systems, the definition of partial transpose needs to be adjusted to accommodate the anticommuting statistical property. There exist two different proposals for fermionic partial transpose and corresponding fermionic negativity, as discussed in Refs. [54, 55] and Refs. [56, 57, 58] respectively. Despite being a “computable entanglement measurement”, fermionic negativity is only analytically tractable in free systems, especially at finite temperatures, and there have been studies based on both the former definition [59, 39, 60, 61] and the latter definition [56, 62, 63, 64, 65]. Therefore, it is desirable to design a quantum Monte Carlo (QMC) algorithm for large-scale simulation of interacting fermionic systems in an unbiased manner, which is the main goal of this letter. Throughout this paper, we adopt the definition in Refs. [56, 57] under which the partial transpose of a Gaussian state remains a Gaussian state. Additionally, instead of utilizing the originally proposed negativity which involves trace norm of partially transposed density matrices (PTDMs) [28], we consider Rényi negativity which involves moments of PTDMs, as done in several previous studies on other systems [32, 33, 45, 37, 40, 39, 47].
In fact, our main result is more broadly applicable. We show that generic PTDMs can be written as a weighted sum of Gaussian states, representing free fermions coupled with auxiliary fields, similar to Grover’s pioneering work on reduced density matrices for EE [66]. Our finding facilitates the calculation of Rényi negativity in a tractable manner, thus establishing it as a powerful tool for characterizing entanglement in mixed states of interacting fermions. We demonstrate this relation using determinant quantum Monte Carlo (DQMC) simulations [67, 68, 69] on two paradigmatic models in the realm of strongly correlated electrons, namely, the Hubbard model and the spinless $t$-$V$ model. These two models on bipartite lattices at half-filling are sign-problem-free and both ground-state and finite-temperature properties can be feasibly simulated within the DQMC framework. The relation between negativity and finite temperature transition in fermionic systems is unveiled.
Partially transposed density matrix in DQMC calculations.—Various definitions of negativity in the literature share a common and central dependency, namely, the partial transpose of the density matrix. In this work, we adopt the partial time-reversal transformation proposed by Shapourian et al. [56, 57] as the fermionic partial transpose.
We begin with the general partitioning of a fermionic lattice model. It is defined using annihilation (creation) operators $c_{j\sigma}^{(\dagger)}$, which satisfy the anticommutation relations $\{c_{j\sigma},c^{\dagger}_{k\sigma^{\prime}}\}=\delta_{jk}\delta_{\sigma\sigma^{\prime}}$, where $j,k=1,\dots,N$ are the labels of the sites and $\sigma,\sigma^{\prime}$ are the indices for internal degrees of freedom such as spin. This lattice system, denoted as $A$, generally exists within a larger space. After tracing out the environment $\bar{A}$, system $A$ typically exists in a mixed state $\rho$. For example, if system $A$ is in contact with a much larger thermal bath at temperature $T$, then we obtain a finite-temperature Gibbs state $\rho=e^{-\beta H}/\mathrm{Tr}e^{-\beta H}$ with $\beta=1/T$ the inverse temperature and $H$ the Hamiltonian of the system $A$. Next, we further divide system $A$ into two parties belonging to two complementary spatial regions respectively, i.e., $A=A_{1}\cup A_{2}$. Then the density matrix acting on Hilbert space $\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ can be expanded as $\rho=\sum_{A_{1},A_{2},A_{1}^{\prime},A_{2}^{\prime}}\rho_{A_{1},A_{2};A_{1}^{\prime},A_{2}^{\prime}}|A_{1}\rangle|A_{2}\rangle\langle A_{1}^{\prime}|\langle A_{2}^{\prime}|$.
The fermionic partial transpose of density matrix $\rho$ with respect to subsystem $A_{2}$, denoted as $\rho^{T_{2}^{f}}$, exhibits a highly succinct mathematical expression in the Majorana basis [56, 57, 70]. Under Majorana basis, an arbitrary density operator can be expressed as a constrained superposition of products of Majorana operators, which are defined as $\gamma_{2j-1,\sigma}=c_{j,\sigma}+c^{\dagger}_{j,\sigma}$ and $\gamma_{2j,\sigma}=-\mathrm{i}(c_{j,\sigma}-c_{j,\sigma}^{\dagger})$. It is found that $\rho^{T^{f}_{2}}$ can be obtained by applying the following transformation to the Majorana operators associated with subsystem $A_{2}$:
$\mathcal{R}_{2}^{f}(\gamma_{j,\sigma})=\text{i}\gamma_{j,\sigma},\quad j\in A_{2}.$ (1)
Remarkably, under this definition, the fermionic partial transpose of a Gaussian state, denoted as $\rho_{0}\sim e^{\frac{1}{4}\bm{\gamma}^{T}W\bm{\gamma}}$, retains its Gaussian nature. The question then pertains to determining the explicit form of $\rho_{0}^{T_{2}^{f}}$ or $W^{T_{2}^{f}}$. To this end, it is important to emphasize that a Gaussian state $\rho_{0}$ can be alternatively characterized by the Green’s function $\Gamma_{kl}=\langle[\gamma_{k},\gamma_{l}]\rangle/2$, which is averaged with respect to $\rho_{0}$ itself and also called covariance matrix. This matrix is connected to the $W$ matrix through the relations $\tanh(-W/2)=\Gamma$ [54, 70]. By employing the definition of $\Gamma$ and the partial transpose in the Majorana basis (refer to Eq. (1)), the partial transpose of the covariance matrix can be formulated as
$\Gamma^{T_{2}^{f}}=\left(\begin{array}[]{cc}\Gamma^{11}&\text{i}\Gamma^{12}\\ \text{i}\Gamma^{21}&-\Gamma^{22}\end{array}\right),$ (2)
where $\Gamma^{bb^{\prime}}$ ($b,b^{\prime}=1,2$) denotes the block comprising the matrix elements with rows pertaining to subsystem $A_{b}$ and columns pertaining to subsystem $A_{b^{\prime}}$. The Gaussian state described by $\Gamma^{T_{2}^{f}}$ precisely yields the $\rho_{0}^{T_{2}^{f}}$ through $\tanh(-{W^{T_{2}^{f}}}/{2})=\Gamma^{T_{2}^{f}}$  [^1] The proof uses the Wick theorem for Majorana monomial, see Ref. [54] and Supplementary Material (SM) [70] for details..
The above discussion in the Majorana basis can be seamlessly transitioned to the complex fermion basis. In complex fermion basis, the Green’s function is defined as $G_{jk}=\langle c_{j}c_{k}^{\dagger}\rangle$, where we have abbreviated the spin indices. Its partially transposed form exhibits also a simple structure
$G^{T_{2}^{f}}=\left(\begin{array}[]{cc}G^{11}&\text{i}G^{12}\\ \text{i}G^{21}&I-G^{22}\end{array}\right),$ (3)
where the superscripts of the blocks $G^{bb^{\prime}}$ indicate the subsystems, akin to the notation of $\Gamma^{bb^{\prime}}$ established earlier. Similar to the Majorana basis, the above Green’s function delineates another Gaussian state which is exactly the partial transpose of the original Gaussian state, i.e., $(\rho_{0}[G])^{T_{2}^{f}}=\rho_{0}[G^{T^{f}_{2}}]$ with $\rho_{0}[G]\sim e^{\mathbf{c}^{\dagger}\ln(G^{-1}-I)\mathbf{c}}$ [78].
It is now pertinent to redirect our attention towards the partial transpose for interacting fermionic systems, whose density matrices are not Gaussian states. Nonetheless, within the framework of DQMC, after Trotter decomposition and Hubbard-Stratonovich decoupling [70], the original two-particle interaction terms are replaced by fermion bilinears coupled with spacetime-dependent auxiliary fields $\mathbf{s}$. The partition function is given by $Z=\sum_{\mathbf{s}}\mathrm{Tr}[\prod_{l=1}^{L_{\tau}}e^{\mathbf{c}^{\dagger}K_{l}[\mathbf{s}]\mathbf{c}}]$, where $L_{\tau}$ is the number of time slices and $\mathbf{c}=(c_{1},\dots,c_{N})^{T}$ (for simplicity we abbreviate spin indices). The density matrix $\rho$ can also be expressed as a weighted sum of Gaussian operators, explicitly $\rho=\sum_{\mathbf{s}}P_{\mathbf{s}}\rho_{\mathbf{s}}$ with $P_{\mathbf{s}}$ representing the weight of configuration $\mathbf{s}$ [66, 70]. After partial transpose, it takes the form
$\rho^{T_{2}^{f}}=\sum_{\bf{s}}P_{\bf{s}}\rho_{\bf{s}}^{T_{2}^{f}},$ (4)
where
$\rho_{{\bf{s}}}^{T_{2}^{f}}=\det\left[G_{{\bf{s}}}^{T_{2}^{f}}\right]\exp\left\{{\mathbf{c}^{\dagger}\ln\left[\left(G_{\mathbf{s}}^{T_{2}^{f}}\right)^{-1}-I\right]\mathbf{c}}\right\}.$ (5)
The aforementioned equations (4) and (5), along with Eq. (3), are the main result of this letter and can be employed to investigate negativity and negativity spectrum within the conventional DQMC framework, fully analogous to the analysis of EE and entanglement spectrum, respectively.
Rényi negativity for Hubbard chain.— We consider the half-filled Hubbard chain with periodic boundary conditions, described by the Hamiltonian
$H=-t\sum_{\langle ij\rangle\sigma}(c^{\dagger}_{i\sigma}c_{j\sigma}+\mathrm{H.c.})+\frac{U}{2}\sum_{i}(n_{i}-1)^{2},$ (6)
which is a sign-problem-free model [69, 70]. We will benchmark DQMC results from two perspectives: (i) a numerical comparison with results obtained from exact diagonalization  [^2] We utilized the definition of fermionic partial transpose in Fock space for exact diagonalization, see the formula Eqs. (S2) and (S3) in SM [70], and (ii) providing a physical explanation for why the negativity is a more competent mixed-state entanglement measurement compared to EE in the context of a quantum-classical crossover [56, 62].
We define the rank-$n$ Rényi negativity as
$\mathcal{E}_{n}=-\frac{1}{n-1}\ln\text{Tr}\left[\left(\rho^{T_{2}^{f}}\right)^{n}\right],$ (7)
where the $n$-th moment of the PTDM, denoted as $\text{Tr}[(\rho^{T_{2}^{f}})^{n}]$  [^3] We note that in the context of studying fermionic systems, an alternative definition for the moments of the PTDM is given by $\text{Tr}(\rho^{T_{2}^{f}}\rho^{T_{2}^{f}\dagger}\cdots)$ [56, 62, 65], which aims to account for the non-Hermitian nature and potential complex eigenvalues associated with the partial time reversal of the density matrix. However, the models under consideration in this letter do not exhibit this issue due to specific symmetries they possess that circumvent the sign problem. , is also referred to as the replica approach of negativity in previous studies [32, 33]. The quantity $\mathcal{E}_{n}$ is formally a direct analog to rank-$n$ Rényi EE $S_{n}(A_{1})=-(\ln\mathrm{Tr}\rho^{n}_{A_{1}})/(n-1)$, where $\rho_{A_{1}}=\mathrm{Tr}_{A_{2}}\rho$ represents the reduced density operator obtained after tracing out subsystem $A_{2}$. Utilizing Eq. (4), we can derive the DQMC expression for measuring, for instance, the rank-two Rényi negativity,
$\mathcal{E}_{2}=$ $-\ln\Big{\{}\sum_{{\bf s}_{1}{\bf s}_{2}}P_{{\bf s}_{1}}P_{{\bf s}_{2}}$ (8)
×$\times\det\left[G_{\bf{s}_{1}}^{T_{2}^{f}}G_{\bf{s}_{2}}^{T_{2}^{f}}+\left(I-G_{\bf{s}_{1}}^{T_{2}^{f}}\right)\left(I-G_{\mathbf{s}_{2}}^{T_{2}^{f}}\right)\right]\Big{\}}$
The distinction between the fermionic partial transpose and the conventional one is presented herein. For bosonic systems, considering that $\mathrm{Tr}[(\rho^{T_{2}})^{2}]=\mathrm{Tr}[\rho^{2}]$, the rank-two Rényi negativity becomes trivial, thereby rendering the minimal meaningful rank as three [32, 33, 47]. However, we show in the occupation number representation that all fermionic PTDMs satisfy ${\rm Tr}[(\rho^{T_{2}^{f}})^{2}]={\rm Tr}[(\rho\hat{X}_{2}(\pi))^{2}]$ with $\hat{X}_{2}\left(\theta\right)=e^{{\rm i}\theta\sum_{j\in A_{2}}{n}_{j}}$ being the disorder operator [70]. Consequently, $\mathcal{E}_{2}$ can reveal the entanglement information of the system. As shown in Fig. 1, the results calculated by DQMC and exact diagonalization show strong agreement in both the zero-temperature and finite-temperature regimes. In the zero-temperature regime, the pattern of the rank-two Rényi negativity exhibits analogous variations to those of the rank-two Rényi EE [66, 81], in response to alterations in the length of subsystem $A_{1}$, denoted as $L_{A_{1}}$. However, at finite temperatures, the negativity maintains a symmetric pattern, which is different from the behavior of EE [82, 81]. As the temperature rises, the magnitude of the negativity increases, resulting in an overall non-zero shift corresponding to a non-zero thermodynamic entropy of $-\ln(\mathrm{Tr}\rho^{2})$.
Figure 1: The variation of the rank-two Rényi negativity $\mathcal{E}_{2}$ for a six-site Hubbard chain with periodic boundary conditions is depicted as a function of the subsystem length $L_{A_{1}}$. The solid lines represent the exact diagonalization results, which agree with the DQMC results at both (a) zero temperature and (b) finite temperatures.
Based on the above observation in finite-temperature regime, we also examine the ratio between $\text{Tr}[(\rho^{T_{2}^{f}})^{n}]$ and $\text{Tr}[\rho^{n}]$ as previous Monte Carlo investigations on bosonic systems [45, 37, 47]
$R_{n}=-\frac{1}{n-1}\ln\left\{\frac{\text{Tr}\left[\left(\rho^{T_{2}^{f}}\right)^{n}\right]}{\text{Tr}[\rho^{n}]}\right\}=\mathcal{E}_{n}-S^{\rm th}_{n},$ (9)
where $S^{\rm th}_{n}=-(\ln\mathrm{Tr}\rho^{n})/(n-1)$ represents the thermodynamic Rényi entropy, which can be computed by considering the subsystem $A_{2}$ as non-existent in $\mathcal{E}_{n}$. The faithful description of mixed-state entanglement necessitates the subtraction of the thermodynamic Rényi entropy $S_{n}^{\rm th}$ from the Rényi negativity $\mathcal{E}_{n}$. In Fig. 2, we display the variations of the negativity ratio and EE with temperature for three distinct lengths, namely $L=6,10,14$. Here, the subsystem $A_{1}$ is selected to be half of the chain, yielding an equal bipartition. As the temperature rises, the EE increases while the negativity ratio asymptotically diminishes to zero for all lengths. This serves as a compelling physical demonstration of the Rényi negativity ratio. In a generic mixed state, both quantum and classical correlations are present, and effective measurement of mixed-state entanglement should exclusively isolate the quantum correlations [28]. In the specific context of finite-temperature Gibbs states, the classical correlation is simply the thermal fluctuations delineated by the thermodynamic entropy $S_{n}^{\rm th}$. Furthermore, at sufficiently low temperatures, the negativity ratio remains constant and establishes a plateau, the length of which is associated with the finite-size gap $1/L$ [62]. As depicted in Fig. 2, it is evident that with an increase in chain length, the plateau becomes narrower. In summary, the monotonic decay of the negativity ratio with rising temperature signifies a crossover from a quantum entangled state to a classical mixed state.
Figure 2: Quantum-classical crossover. The scaled Rényi negativity ratio $R_{2}/L$ and EE $S_{2}/L$ of the half-filled Hubbard chain under a half-chain bipartition vary as functions of temperature. As the temperature rises, the scaled EE for different lengths increases and converges, indicating a dominance of volume law at high temperatures. Meanwhile, the negativity ratio begins to vanish once the temperature reaches a critical value associated with the finite-size gap $1/L$ [62].
Finite-temperature transition in $t$-$V$ model.—To demonstrate the efficacy of Rényi negativity ratio in detecting finite-temperature phase transition, we further consider the half-filled spinless $t$-$V$ model on a square lattice with periodic boundary conditions [83, 84, 82],
$H=-t\sum_{\langle i,j\rangle}(c_{i}^{\dagger}c_{j}+c_{j}^{\dagger}c_{i})+V\sum_{\langle i,j\rangle}\left(n_{i}-\frac{1}{2}\right)\left(n_{j}-\frac{1}{2}\right),$ (10)
where both the hopping and the interaction involve only nearest neighbors. In the presence of a finite coupling parameter $V$, this model exhibits a charge density wave (CDW) ground state and undergoes a phase transition from the CDW phase to a metallic phase at finite temperature, with critical behavior falling within the 2D Ising universality class [84, 85]. In the following, we focus on a specific coupling strength, $V/t=2$, where the critical temperature is estimated to be approximately $T_{c}/t\approx 1.0$, a value lower than the 2D Ising result of $1.120.56|V/t|=$ 1.12 [84].
This model is also a sign-problem-free model [86, 87, 88, 89]. However, for models with larger dimensions or stronger interaction strengths, the direct sampling of Rényi negativity using Eq. (8) becomes inaccurate and challenging to converge, as a result of the occurrence of spikes [90] or the non-Gaussian distribution of Grover determinants $\det g_{x}=\det[G_{\bf{s}_{1}}^{T_{2}^{f}}G_{\bf{s}_{2}}^{T_{2}^{f}}+(I-G_{\bf{s}_{1}}^{T_{2}^{f}})(I-G_{\mathbf{s}_{2}}^{T_{2}^{f}})]$ [91]. We implement an incremental algorithm for rank-$n$ Rényi negativity, analogous to the controllable incremental algorithm for Rényi EE [92, 93, 91, 94], the spirit of which is to measure $(\det g_{x})^{1/N_{\rm inc}}$ instead of $(\det g_{x})$ to circumvent the sampling of an exponentially small quantity with exponentially large variance [95]. It is important to note that there is a sign ambiguity in the $N_{\rm inc}$-th root. In the Supplementary Material [70], we prove that the Grover determinant $\det g_{x}$ is always real and non-negative for two classes of sign-free models, represented by the Hubbard model and the spinless $t$-$V$ model, respectively.
Figure 3: The finite temperature transition in the spinless $t$-$V$ model is detected by the area-law coefficient of the Rényi negativity ratio as a function of temperature. The geometry of the bipartition is illustrated in the inset on the right. A vertical arrow, colored in gray, indicates the position of the shared peak, with half of it aligning with the transition point determined in previous studies [84, 82]. The left inset shows the linear scaling of the area-law coefficient at the critical point with $\ln L$.
As illustrated in the right inset of Fig. 3, we designate the lower left corner with dimensions $(L/2)\times(L/2)$ as subsystem $A_{1}$, resulting in an area-law coefficient of the Rényi negativity ratio of $R_{2}/L$. The main plot of Fig. 3 presents $R_{2}/L$ as a function of temperature for various system sizes, demonstrating a notably distinct finite-size characteristic compared to the intersection of mutual information [96, 97, 98, 82]. Remarkably, unlike the Hubbard model in Fig. 2 or the previous study on the $2+1\mathrm{D}$ transverse field Ising model [47], the Rényi negativity ratio does not exhibit a monotonic decrease with rising temperature. Instead, for varying lattice sizes, a shared local maximum appears at approximately twice the transition temperature, $2T_{c}/t\approx 2.1$. The inclusion of the prefactor $2$ aligns with the rank of the Rényi negativity ratio under consideration, consistent with earlier discussion on the critical behavior of Rényi negativity [51, 53, 47]. The peak of $R_{2}/L$ exhibits a logarithmic divergence with system size $L$ as shown in the left inset of Fig. 3, while the area law is well preserved far away from the critical point  [^4] This beyond-area-law scaling around the finite temperature critical point is also observed for other values of $V$, a different bipartite geometry, and a different lattice. Refer to the SM for additional complementary plots [70].. Based on symmetry considerations, it was argued that the entanglement negativity inherits the singularity of the specific heat at a finite temperature transition [53, 47], and for the 2D Ising transition, the specific heat has a logarithmic divergence. However, in our fermionic scenario, the quantity showing this divergence is $R_{2}/L$ rather than its temperature derivative. The underlying cause of the logarithmic divergence of $R_{2}/L$ at $2T_{c}$ thus remains an unresolved issue.
Conclusions and outlook.—We showed that the partially transposed density matrix for interacting fermions, akin to the reduced density matrix, can be expanded as a weighted sum of Gaussian states representing free fermions, thereby paving the way for the study of mixed-state entanglement in strongly correlated fermionic systems. This main result was employed to implement an algorithm to compute the rank-$n$ Rényi negativity for interacting fermionic systems within the DQMC framework. We presented the first study of the rank-two Rényi negativity for the half-filled Hubbard chain and the spinless $t$-$V$ model on a square lattice. Remarkably, we found that the area law coefficient of the negativity ratio exhibits a logarithmic singular peak at twice the finite-temperature transition point for all lattice sizes under consideration.
There are several potential future research directions to consider. The first direction is to investigate the finite-temperature entanglement of various interacting fermionic models, especially those with transition points that belong to different universality classes, such as the 3D Hubbard model which owns a transition belonging to $\mathrm{O}(3)$ universality class [100]. Further, exploring the entanglement in other types of mixed states, such as tripartite ground states of topological [57] and gapless systems [39], and measurement-induced mixed states [101, 102] presents an intriguing avenue for further research. Next, exploring the finite-size scaling laws of negativity in interacting fermionic systems, especially the beyond-area-law behaviors that reveal universal properties of quantum systems, could also be intriguing. In particular, our finding of the $L\ln L$ scaling of the negativity ratio at the critical point in the spinless $t$-$V$ model may indicate long-range entanglement contribution [47, 65], which warrants further investigation. Moreover, our results are applicable to the continuous-time QMC method, offering an opportunity to study the mixed-state entanglement of realistic correlated materials through combining with dynamical mean-field theory [82, 103, 104]. For example, to fit into the hybridization expansion algorithm, the bath can be first traced out [78, 66] and then the density matrix of the impurity can be obtained using the Green’s functions and density correlation functions [65, 105]. Another interesting example is to integrate the rank-two Rényi negativity into the interaction expansion algorithm by making use the identity ${\rm Tr}[(\rho^{T_{2}^{f}})^{2}]={\rm Tr}[(\rho X_{2}(\pi))^{2}]$, where the disorder operator, expressed in the interaction picture as $X_{2}(\tau)\propto\prod_{j\in A_{2},\sigma}(n_{i\sigma}(\tau)-\frac{1}{2})$, is nothing just some additional interaction vertices only at $A_{2}$ subregion. Finally, another proposal for realistic materials is to integrate the Rényi negativity into the constrained-path auxiliary-field QMC method [106, 107, 108, 109] which controls the sign problem.
Acknowledgements.
Acknowledgements.—We thank Tarun Grover for helpful discussions and comments on the draft. This work is supported by the National Key R&D Program of China (Grant No. 2022YFA1402702, No. 2021YFA1401400), the National Natural Science Foundation of China (Grants No. 12274289), the Innovation Program for Quantum Science and Technology (under Grant No. 2021ZD0301902), Yangyang Development Fund, and startup funds from SJTU. The computations in this paper were run on the Siyuan-1 and $\pi$ 2.0 clusters supported by the Center for High Performance Computing at Shanghai Jiao Tong University.